### v1

@@ Abstract
We here present an improved version of the Sortition Foundation’s GroupSelect software package, which aims to repeatedly allocate participants of a deliberative process to discussion groups in a way that balances demographics in each group and maximises distinct meetings over time. Our result, GroupOpt, significantly outperforms the prior algorithmic approach Legacy. We also add functionalities to the GroupSelect software to help the end user. The GroupOpt algorithm utilises random shuffles and Pareto swaps to find a locally optimal solution that maximises demographic balance and minimises the number of pairwise previous meetings, with the relative importance of these two metrics defined by the user.
@@ 1 Introduction
Theoretical attempts to improve upon Legacy introduced in Barrett et al [10] rely on the cutting-edge Gurobi optimisation solver, which is free for academic licenses but prohibitively expensive for industry applications. [^1] https://www.gurobi.com/request-a-price-quote/ (accessed 22/07/2024). We are aware of a non-profit Gurobi track for non-profit use is for “qualifying non-profit organisations working to support at least one of the United Nations’ Sustainable Development Goals”. However, we see this as insufficient for our purposes as we cannot guarantee that all end-users would satisfy this criteria, particularly in light of their clarification that a “government agency requiring Gurobi for operations” does not count as non-profit, and government agencies are significant drivers of large-scale CAs While open-source solvers have improved in recent years, they still fail on realistic table allocation use cases due to the exponential parameter growth [11]. Therefore, we require a heuristic approach that leverages the insights from an optimised approach and outperforms the current art.
In this work we introduce a way to achieve the goals of a good allocation algorithm: (1) maximal unique pairs, (2) balanced demographics, and (3) a clustering option. In previous investigations of the Legacy algorithm [10], it was found that the approach struggled to satisfy objective (1), particularly over large numbers of rounds. This was because the heuristic approach is effectively deterministic up to swaps within sets of demographically identical individuals. Furthermore, the Legacy algorithm only considers first meetings between pairs, with no preference over subsequent meetings: so a pair meeting 10 times is considered equally as valuable to a process as a pair meeting twice.
Our new approach utilises random shuffles and local Pareto optimisation to attain a stronger meeting profile than that achieved by Legacy  [12]. The updated version of the GroupSelect app, containing the improved heuristic algorithm GroupOpt, is available here.
@@ 2 Notation
We define a setting of I agents in the deliberative panel, J discussion groups (henceforth referred to as ‘tables’, even if they represent an online space), and $K$ rounds of discussion. In each round $k\in K$, each agent $i\in\textbf{I}$ is allocated to table $j\in\textbf{J}$, using an allocation function $A_{k}:\textbf{I}\rightarrow\textbf{J}$. Further notation is specific to the GroupOpt approach:
@@ 3 Methodology and Heuristic Algorithm
The GroupOpt algorithm is described in Algorithm 1:
Process 1 GroupOpt Algorithm
@@
4:     Randomly allocate clustered agents on a subset of $\mid\textbf{J}\mid^{c}$ cluster tables
@@
For example, for binary demographic $d=$age and value $v_{d}=$young, if the age split is 50/50 in the panel and we are on a table with 2 young and 8 old, then $p_{v_{d},k}^{j}=0.2$ and $\Delta_{j,d,k}=0.6$ (the distance from optimal for old + young). We can swap one of the 8 old people for anyone, as both options do not Pareto harm the table for age, or a young person for a young person, but swapping a young person for an old person would push us further from 50/50 so would not be allowed. Therefore, $P(i,i^{\prime},j,d,k)=1$ if $i$ is old and $i^{\prime}$ is young, $P(i,i^{\prime},j,d,k)=0$ if $d(i)=d(i^{\prime})$, and $P(i,i^{\prime},j,d,k)=-1$ if $i=$young, $i^{\prime}=$old.
@@ 3.1 Algorithmic Implementation
The code implementation of GroupOpt is packaged with an updated version of the GroupSelect interface with some additional error messages and user guidance around input variables. In particular, interface improvements are:
@@
  • A more accurate calculation of the maximum possible number of links based on the size of the panel and tables, as shown in Barrett et al [10]. The previous interface gave a naive maximum calculation based on the combinatoric number of pairs in the panel.
  • Clustering input updates: as clustering is mainly used to cluster correlated variables (namely consenting to media and conversational tracking), the option to cluster these variables separately is inefficient and limits Legacy’s performance in other areas. The algorithm ensures that all clustering constraints are satisfied before moving on to diversification considerations. As the variables are correlated, it makes more sense to collect the clustering constraints into a single binary variable, where a single value in the variable denotes that individuals with this value should be clustered together.
@@ 3.2 Experiment Set-up
For each experiment with a clustering variable, in GroupOpt, we used a binding number of cluster tables as this is most closely comparable to Legacy. For example, if tables are size 5 and there are 15 cluster agents, 3 tables will be dedicated exclusively to clustering. This means that no cluster agents will ever meet non-cluster agents, severely limiting the number of meetings achievable; even extending to 4 tables would be a significant relaxation. This also limits representation; for example, if all 15 cluster agents are also male, then balance on the cluster tables will be artificially low; the new GroupSelect interface would normally suggest a larger number of cluster tables in this scenario. While this approach allows for valid comparisons to Legacy, experiments represent a lower bound on the possible performance in terms of both meetings and demographics in experiments involving clustering (3/7ths of all experiments). We also chose $n_{S}=5$ rounds of swaps: in the most complex case, this resolved in a runtime of $<$10 seconds, so better performance is likely possible with further rounds of swaps. Finally, GroupOpt implementations used a flat $p_{s}=0.5$ parameter for mixing between meetings and demographic prioritisation. A different value could improve meetings at the expense of diversity, or vice versa, even more than the below: so again, results are a lower bound on the possible performance for either meetings or demographics.
We note that the our experimental variables range from binary (two possible values) to five possible values. It has been observed in previous implementations of Legacy that the algorithm has struggled to allocate data sets with many-option variables, for example participants’ granular states or cantons. We truncate our experiments at five options for two reasons. The maximum is as high as five to show that the new approach can handle variables with more than two or three options. The maximum is as low as five because more options than this does not make sense as input for the algorithm: if a variable has levels that are granular and only held by a small subset of participants, it is objectively difficult to fairly allocate these between tables (i.e. how should the algorithm allocate the 3 participants from city X across 10 tables), and we recommend such variables are pre-processed into broader groups (e.g. region). The GroupSelect interface is being updated to automatically suggest where and how variables levels should be clustered before running the algorithm.
@@ 3.3 Evaluating the Algorithm
  • Meetings: influenced by the results in Barrett et al [10], we use a geometric saturation function $f$ to evaluate algorithmic performance. The function $f$ evaluates the value to the process of the number of times each pair meets. We wish this function to be increasing (i.e. meetings add value) and concave (i.e. a pair’s second meeting is less valuable than their first, as they have already exchanged ideas). Among the candidates for a ‘good’ saturation function, there are two promising families. The first is harmonic, where each pair’s $n^{th}$ meeting generates a marginal utility of $1/n$. This leads to a slow decrease in marginal utility, placing significant value on later meetings. However, it leads to strange comparative behaviour: two pairs meeting for a second time ($/2$ 2) generate the same utility as a single pair meeting for the first time (1), but two pairs meeting for the third time ($2/3$) generate significantly more utility than a pair meeting for the second time ($1/2$). We therefore turn to the geometric saturation function, where each pair’s $n^{th}$ meeting generates marginal utility of $a^{n}$ for some constant $10<a<$ 1 (we use $a=0.5$). This leads to a faster decrease in marginal utility than the harmonic function, but has the appealing ‘self-similarity’ property whereby once everyone has met $x$ times, the problem effectively resets (as the $x+1^{st}$ meeting is worth precisely $a$ $x^{th}$ meetings, which is the same as a second meeting being worth $a$ first meetings). In our evaluation, each pair’s total meeting score over the process is aggregated to get a direct comparison of meeting qualities between Legacy and GroupOpt
@@ 4 Results
We present the set-up and results from 66 experiments, each run for both Legacy and GroupOpt.
@@ 4.1 Experiment Outcomes
For every single experiment, the meeting score is higher under GroupOpt than Legacy. This difference becomes more significant for longer runs, as shown in Table 4.1.
@@
Thus, performance improves at a higher rate when the algorithm runs for longer time periods. We also note that, aside from the artificially constrained $\mid\textbf{I}\mid=120$ case with large groups, more participants with a fixed table size also leads to a greater performance increase for GroupOpt over Legacy. To give an illustration of how this works, we can focus on a specific experiment for the $\mid\textbf{I}\mid=100$, $K=10$ rounds example in Figures 1(a) and 1(b).
@@
Following the approach in Barrett et al [10], we can also see how our GroupOpt algorithm compares to Legacy when we judge it by its own performance criteria, maximising the number of first meetings. Based on previous calculations of the minimum number of repeated meetings between any two rounds, we can calculate the maximum number of first meetings that are possible, given the panel size and number of tables [^3] Note that in many cases this number will not be achievable, as reallocating in a way that generates the minimum number of repeated meetings may only be possible once. Under this approach, Legacy achieves 62.6% of possible first meetings averaged over non-clustering implementations. GroupOpt achieves 76.9 of all possible first meetings.
When considering optimal performance metric $excess$, GroupOpt outperforms Legacy on all but one of the experiments. The exception is a simple case with $\textbf{I}=30$, $\textbf{J}=2$, $K=3$, where GroupOpt leads to three fewer pairs meeting than Legacy (a proportionate difference of 0.7% of all pairs). Across all experiments, the average improvement generated by GroupOpt (i.e. $excess^{GS}-excess^{GB}$) is 12.7%.
Differences in average diversity scores $\bar{\Delta}^{A,\textsc{Legacy}{}}-\bar{\Delta}^{A,\textsc{Group\-Opt}{}}$ are shown in Figure 2:
Figure 2: Histogram of experiments in which GroupOpt attains better (positive values), the same (0 values) or worse (negative values) diversity scores than Legacy
@@
Finally, we note that GroupOpt is trivially better than random allocation by construction, as the initial allocation is based on randomness and the set of Pareto swaps explicitly improve upon this.
@@ 4.2 Comparisons with Optimal
The attentive reader may note that we have omitted direct comparisons with the algorithm from Barrett et al [10]. This is because their approach only seeks optimal allocations on a round-by-round basis, and therefore does not represent the best possible performance. Having established that GroupOpt outperforms Legacy, in this Section we explore how close the algorithm gets to true optimum.
@@
The excess number of pairs who do not meet under GroupOpt (Figure 3) ranges from a minimum of 0.5% of all pairs to a maximum of 24.6%, with an average score of 9.5%. The cluster of results with an excess of under 10% (59.0% of experiments) indicates that the algorithm performs close to optimal in most settings, while the relatively low maximum value indicates that even in complex situations with multiple diversification variables, performance does not leak to a dramatic extent. For context, the worst-performing experiment for Legacy sees 74.0% of pairs not meeting who theoretically could.
Figure 3: $excess$ values for GroupOpt experiments: the proportion of pairs that don’t meet over the panel but theoretically could
@@ 5 Future Developments
Given the highly promising experimental results, we are currently working on implementing the GroupOpt algorithm and an updated GroupSelect interface with the Sortition Foundation. Beyond implementation, these results leave open questions. The persistence of some excess meetings over an optimal performance is partially a function of the problem setting: the possibility of achieving the minimum $L_{R}^{*}(k,k^{\prime})$ repeated meetings between two rounds does not necessarily imply that this result scales for 3+ rounds, as the problem becomes exponentially more complex with more rounds. However, the higher $excess$ in the worse-performing experiments indicates that some further improvement on GroupOpt is likely attainable. Additionally, a user study on how practitioners deploy this algorithm in the field is vital to ensuring that both the algorithm is used as intended and the front-end features, such as app design and layout, are optimised for users’ convenience.

### v2

@@ Abstract
We here present an improved version of the Sortition Foundation’s GroupSelect software package, which aims to repeatedly allocate participants of a deliberative process to discussion groups in a way that balances demographics in each group and maximises distinct meetings over time. Our result, Dream, significantly outperforms the prior algorithmic approach Legacy. We also add functionalities to the GroupSelect software to help the end user. The Dream algorithm utilises random shuffles and Pareto swaps to find a locally optimal solution that maximises demographic balance and minimises the number of pairwise previous meetings, with the relative importance of these two metrics defined by the user.
@@ 1 Introduction
Theoretical attempts to improve upon Legacy introduced in Barrett et al [10] rely on the cutting-edge Gurobi optimisation solver, which is free for academic licenses but prohibitively expensive for industry applications. [^1] https://www.gurobi.com/request-a-price-quote/ (accessed 22/07/2024). We are aware of a non-profit Gurobi track for non-profit use, which is for “qualifying non-profit organisations working to support at least one of the United Nations’ Sustainable Development Goals”. However, we see this as insufficient for our purposes as we cannot guarantee that all end-users would satisfy this criteria, particularly in light of their clarification that a “government agency requiring Gurobi for operations” does not count as non-profit, and government agencies are significant drivers of large-scale CAs While open-source solvers have improved in recent years, they still fail on realistic table allocation use cases due to the exponential parameter growth [11]. Therefore, we require a heuristic approach that leverages the insights from an optimised approach and outperforms the current art.
In this work we introduce a way to achieve the above goals of a good allocation algorithm: (1) balanced demographics, (2) a clustering option, and (3) maximal unique pairs. In previous investigations of the Legacy algorithm [10], it was found that the approach struggled to satisfy objective (3), particularly over large numbers of rounds. This was because the heuristic approach is effectively deterministic up to swaps within sets of demographically identical individuals. Furthermore, the Legacy algorithm only considers first meetings between pairs, with no preference over subsequent meetings: so a pair meeting 10 times is considered equally as valuable to a process as a pair meeting twice.
Our new approach utilises random shuffles and local Pareto optimisation to attain a stronger meeting profile than that achieved by Legacy  [12]. The updated version of the GroupSelect app, containing the improved heuristic algorithm Dream (Diverse Representation and Engagement Allocation Method), is available here.
@@ 2 Notation
We define a setting of I agents in the deliberative panel, J discussion groups (henceforth referred to as ‘tables’, even if they represent an online space), and $K$ rounds of discussion. In each round $k\in K$, each agent $i\in\textbf{I}$ is allocated to table $j\in\textbf{J}$, using an allocation function $A_{k}:\textbf{I}\rightarrow\textbf{J}$. Further notation is specific to the Dream approach:
@@
  • A subset of tables $\textbf{J}^{c}$ are reserved for individuals with clustering requirements
@@ 3 Methodology and Heuristic Algorithm
The Dream algorithm is described in Algorithm 1:
Process 1 Dream Algorithm
@@
4:     Randomly allocate clustered agents on a subset of cluster tables $\textbf{J}^{c}$
@@
For example, for binary demographic $d=$age and value $v_{d}=$young, if the age split is 50/50 in the panel and we are on a table with 2 young and 8 old, then $p_{v_{d},k}^{j}=0.2$ and $\Delta_{j,d,k}=0.6$ (the distance from optimal for old + young). We can swap one of the 8 old people for anyone, as both options do not Pareto harm the table for age, or a young person for a young person, but swapping a young person for an old person would push us further from 50/50 so would not be allowed. Therefore, $P(i,i^{\prime},j,d,k)=1$ if $d(i)=$ old and $d(i^{\prime})=$ is young, $P(i,i^{\prime},j,d,k)=0$ if $d(i)=d(i^{\prime})$, and $P(i,i^{\prime},j,d,k)=-1$ if $d(i)=$young, $(i^{\prime})=$old.
@@ 3.1 Algorithmic Implementation
The code implementation of Dream is packaged with an updated version of the GroupSelect interface with some additional error messages and user guidance around input variables. In particular, interface improvements are:
@@
  • A more accurate calculation of the maximum possible number of links based on the size of the panel and tables, as shown in Barrett et al [10]. The previous interface gave a naïve maximum calculation based on the combinatoric number of pairs in the panel.
  • Clustering input updates: as clustering is mainly used to cluster correlated variables (namely consenting to media and conversational tracking), the option to cluster these variables separately is inefficient and limits Legacy’s performance in other areas. It makes more sense to collect the clustering constraints into a single binary variable, where a single value in the variable denotes that individuals with this value should be clustered together. The algorithm ensures that all clustering constraints are satisfied before moving on to diversification considerations.
@@ 3.2 Experiment Set-up
For each experiment with a clustering variable, in Dream, we used a binding number of cluster tables as this is most closely comparable to Legacy. For example, if tables are size 5 and there are 15 cluster agents, 3 tables will be dedicated exclusively to clustering. This means that no cluster agents will ever meet non-cluster agents, severely limiting the number of meetings achievable; even extending to 4 tables would be a significant relaxation. This also limits representation; for example, if all 15 cluster agents are also male, then balance on the cluster tables will be artificially low; the new GroupSelect interface would normally suggest a larger number of cluster tables in this scenario. While this approach allows for valid comparisons to Legacy, experiments represent a lower bound on the possible performance in terms of both meetings and demographics in experiments involving clustering (3/7ths of all experiments). We also chose $n_{S}=5$ rounds of swaps: in the most complex case, this resolved in a runtime of $<$10 seconds, so better performance is likely possible with further rounds of swaps. Finally, Dream implementations used a flat $p_{s}=0.5$ parameter for mixing between meetings and demographic prioritisation. A different value could improve meetings at the expense of diversity, or vice versa, even more than the below: so again, results are a lower bound on the possible performance for either meetings or demographics.
We note that the our experimental variables range from binary (two possible values) to five possible values. It has been observed in previous implementations of Legacy that the algorithm has struggled to allocate data sets with many-option variables, for example participants’ granular states or cantons. We truncate our experiments at five options for two reasons. The maximum is as high as five to show that the new approach can handle variables with more than two or three options. The maximum is as low as five because more options than this does not make sense as input for the algorithm: if a variable has levels that are granular and only held by a small subset of participants, it is objectively difficult to fairly allocate these between tables (i.e. how should the algorithm allocate the 3 participants from city X across 10 tables). We recommend such variables are pre-processed into broader groups (e.g. region). The GroupSelect interface is being updated to automatically suggest where and how variables levels should be clustered before running the algorithm.
@@ 3.3 Evaluating the Algorithm
  • Meetings: influenced by the results in Barrett et al [10], we use a geometric saturation function $f$ to evaluate algorithmic performance. The function $f$ evaluates the value to the process of the number of times each pair meets. We wish this function to be increasing (i.e. meetings add value) and concave (i.e. a pair’s second meeting is less valuable than their first, as they have already exchanged ideas). Among the candidates for a ‘good’ saturation function, there are two promising families. The first is harmonic, where each pair’s $n^{th}$ meeting generates a marginal utility of $1/n$. This leads to a slow decrease in marginal utility, placing significant value on later meetings. However, it leads to strange comparative behaviour: two pairs meeting for a second time ($/2$ 2) generate the same utility as a single pair meeting for the first time (1), but two pairs meeting for the third time ($2/3$) generate significantly more utility than a pair meeting for the second time ($1/2$). We therefore turn to the geometric saturation function, where each pair’s $n^{th}$ meeting generates marginal utility of $a^{n}$ for some constant $10<a<$ 1 (we use $a=0.5$). This leads to a faster decrease in marginal utility than the harmonic function, but has the appealing ‘self-similarity’ property whereby once everyone has met $x$ times, the problem effectively resets (as the $x+1^{st}$ meeting is worth precisely $a$ $x^{th}$ meetings, which is the same as a second meeting being worth $a$ first meetings). In our evaluation, each pair’s total meeting score over the process is aggregated to get a direct comparison of meeting qualities between Legacy and Dream
@@ 4 Results
We present the set-up and results from 66 experiments, each run for both Legacy and Dream.
@@ 4.1 Experiment Outcomes
For every single experiment, the meeting score is higher under Dream than Legacy. This difference becomes more significant for longer runs, as shown in Table 4.1.
@@
Thus, performance improves at a higher rate when the algorithm runs for longer time periods. We also note that, aside from the artificially constrained $\mid\textbf{I}\mid=120$ case with large groups, more participants with a fixed table size also leads to a greater performance increase for Dream over Legacy. To give an illustration of how this works, we can focus on a specific experiment for the $\mid\textbf{I}\mid=100$, $K=10$ rounds example in Figures 1(a) and 1(b).
@@
Following the approach in Barrett et al [10], we can also see how our Dream algorithm compares to Legacy when we judge it by its own performance criteria: maximising the number of first meetings. Based on previous calculations of the minimum possible number of repeated meetings between any two rounds, we can calculate the maximum number of first meetings that are possible, given the panel size and number of tables. [^3] Note that in many cases this number will not be achievable, as reallocating in a way that generates the minimum number of repeated meetings may only be possible once Under this approach, Legacy achieves 62.6% of possible first meetings averaged over non-clustering implementations. Dream achieves 76.9% of all possible first meetings.
When considering optimal performance metric $excess$, Dream outperforms Legacy on all but one of the experiments. The exception is a simple case with $\textbf{I}=30$, $\textbf{J}=2$, $K=3$, where Dream leads to three fewer pairs meeting than Legacy (a proportionate difference of 0.7% of all pairs). Across all experiments, the average improvement generated by Dream (i.e. $excess^{\textsc{Legacy}{}}-excess^{\textsc{Dream}{}}$) is 12.7%.
Differences in average diversity scores $\bar{\Delta}^{A,\textsc{Legacy}{}}-\bar{\Delta}^{A,\textsc{Dream}{}}$ are shown in Figure 2:
Figure 2: Histogram of experiments in which Dream attains better (positive values), the same (0 values) or worse (negative values) diversity scores than Legacy
@@
Finally, we note that Dream is trivially better than random allocation by construction, as the initial allocation is based on randomness and the set of Pareto swaps explicitly improve upon this.
@@ 4.2 Comparisons with Optimal
The attentive reader may note that we have omitted direct comparisons with the algorithm from Barrett et al [10]. This is because their approach only seeks optimal allocations on a greedy round-by-round basis, and therefore does not represent the best possible performance. Having established that Dream outperforms Legacy, in this Section we explore how close the algorithm gets to true optimum.
@@
The excess number of pairs who do not meet under Dream (Figure 3) ranges from a minimum of 0.5% of all pairs to a maximum of 24.6%, with an average score of 9.5%. The cluster of results with an excess of under 10% (59.0% of experiments) indicates that the algorithm performs close to optimal in most settings, while the relatively low maximum value indicates that even in complex situations with multiple diversification variables, performance does not leak to a dramatic extent. For context, the worst-performing experiment for Legacy sees 74.0% of pairs not meeting who theoretically could.
Figure 3: $excess$ values for Dream experiments: the proportion of pairs that don’t meet over the panel but theoretically could
@@ 5 Future Developments
Given the highly promising experimental results, we are currently working on implementing the Dream algorithm and an updated GroupSelect interface with the Sortition Foundation. Beyond implementation, these results leave open questions. The persistence of some excess meetings over an optimal performance is partially a function of the problem setting: the possibility of achieving the minimum $L_{R}^{*}(k,k^{\prime})$ repeated meetings between two rounds does not necessarily imply that this result scales for 3+ rounds, as the problem becomes exponentially more complex with more rounds. However, the higher $excess$ in the worse-performing experiments indicates that some further improvement on Dream is likely attainable. Additionally, a user study on how practitioners deploy this algorithm in the field is vital to ensuring that both the algorithm is used as intended and the front-end features, such as app design and layout, are optimised for users’ convenience.